In a straight line motion the distance travelled is propertional to the square root of the time taken. The acceleration of the particle is proportional to

Grade:11

1 Answers

Mohit Dhaka
7 years ago
S=t^1/2 so differentiating it v=1/2t^1/2 once more differentiating it a=-1/4t^3/2 so accn is inversely proportional to square root of cube of t =t^3/2